OpenStudy (anonymous):

A 3 ft by 5 ft wooden board has three equally sized circles cut out of it. A penny will be randomly tossed onto the board. What is the probability the penny will fall through any of the holes? @david111 @acxbox22

OpenStudy (anonymous):

I got 16%

OpenStudy (acxbox22):

dont you have measurements for the circles?

OpenStudy (anonymous):

1 inch circles .

OpenStudy (acxbox22):

1 inch in area or radius or diameter?

OpenStudy (acxbox22):

and how did you get 16%?

OpenStudy (anonymous):

there is a 5.29% chance it falls in the center hole so I added 5.29% 3 times for the three holes and got 16% and it is a 1 inch diameter.

OpenStudy (anonymous):

area of the board = 3*5 = 15 sq. ft. area of a circle = pi*R^2

OpenStudy (acxbox22):

find the area of the circle and multiply by 3 find area of rectangle circle total area/rectangle area is your probability

OpenStudy (campbell_st):

convert the board measurements to inches... before finding the area... then the probability will be P(penny falls through) = area of 3 holes/area of the board

OpenStudy (anonymous):

i got 8.3%

OpenStudy (campbell_st):

@Faith_Rochelle do you think you could post the entire question.... as it seems you are given some information about whether the penny falls though or not

OpenStudy (acxbox22):

hint: area of one circle is about 0.785 area of rectangle is 2160

OpenStudy (acxbox22):

answer is about 0.011%

OpenStudy (anonymous):

faith what is choices

OpenStudy (anonymous):

@campbell_st that is the entire question . @acxbox22 that is not a choice. 3% 16% 32% 63% ^^^ choices .

OpenStudy (campbell_st):

so how do you know 5.29% of the centre circle...?

OpenStudy (acxbox22):

i know how you got %16 percent and it is right the question you gave us is incomplete the circles are probably 1 feet diameter or the board is 3x5inches

OpenStudy (anonymous):

thats 1 foot u have 1 inch diamter

OpenStudy (acxbox22):

exactly..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I meant one foot . I am sooo sorry . I am just really sleepy . My bad .

OpenStudy (acxbox22):

yeah your answer is 16%

OpenStudy (anonymous):

Mrs. Brown gives two tests to her class. 30% of the class passed both tests and 45% of the class passed the first test. What percent of those who passed the first test also passed the second test?

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

I asked my friend that is a senior about that one ^^ I just posted and she said it is 75% because you add 30 and 45 .. but that does not seem right to me . Any suggestions ? @acxbox22

OpenStudy (anonymous):

66.67%

OpenStudy (anonymous):

In other words 67% ? @david111 & how did you get that ?

OpenStudy (acxbox22):

the answer is in your question! it is 30%

OpenStudy (campbell_st):

is the daimeter 1 ft...?

OpenStudy (acxbox22):

your question asks how many people passed the first and second test and it say 30% passed both tests!

OpenStudy (acxbox22):

troll question

OpenStudy (anonymous):

The choices are 15% 33% 67% 75%

OpenStudy (acxbox22):

according to my logic of course

OpenStudy (anonymous):

it has to be 67%

OpenStudy (acxbox22):

@david111 will you please use your own mind and not rely on answers from yahoo and if you believe that is the answer than will you please explain

OpenStudy (anonymous):

yea its 75%

OpenStudy (anonymous):

30+45% passed test = 75%

OpenStudy (anonymous):

actually 55% didnt pass second test

OpenStudy (acxbox22):

well you cant just throw around percents like that they are not regular integers

OpenStudy (anonymous):

i honostly think its 15% campbell_st

OpenStudy (anonymous):

ignore the campbell